OpenStudy (anonymous):

i need help with volumes of prisms and cylinders (8 problems)(medal)

Directrix (directrix):

@desfields901 Tip: Post one problem per thread.

OpenStudy (anonymous):

i have only 2 or 3 and alright

Directrix (directrix):

As before, V = pi * r^2 * h 271.4 = pi * 6^2 * h 271.4 = 3.14 * 6^2 * h 271.4 = 3.14 * 36 * h @desfields901 What is the product of 3.14 and 36 = ?

Directrix (directrix):

Here is where we are: 271.4 = ( 3.14 * 36) * h --> getting ready to solve for h in the next step.

Directrix (directrix):

Multiply---> 3.14 times 36 so we can finish this, please.